OpenStudy (anonymous):

use the distributive property to simplify the expression below -5x+15/-5

OpenStudy (waleed_imtiaz):

Distributive property ? how .? it is.... a(b+c)=ab + ac..... just simplifying will give -5x-3=0 -5x=3 ; x=-3/5

OpenStudy (anonymous):

 this is what the problem looks like